What is the pH of a solution that consists of 0.50 M H2C6H6O6 (ascorbic acid) and 0.75 M NaHC6H6O6 (sodium ascorbate)? What is the pH if 20.0 mL of a 0.280 M perchloric acid is added to 200.mL of the ascorbic acid sodium ascorbate solution? What is the pH if 30.0 mL of a 0.125 M sodium hydroxide is added to 200.mL of the ascorbic acid sodium ascorbate solution? The ionization constant for ascorbic acid is 6.8 x 10–5
Answer #1

pH of ascorbic acid and sodium ascorbate solution = 4.35

pH after addition of perchloric acid = 4.306

pH after addition of sodium hydroxide = 4.373

0.5 m ascorbic acid and 0.75 M Sodium ascorbate 108ms a bubbex Solution ka = 6.8*105 pka = -log ka = -log (6.8x10-5) [pka = 4Abtex addition of a perchloric acid millimoles of ascorbic acid : 100 + x = 100+5.6 = 105.6 mmol millimoles of sodium ascorbaAfter addition 06 Nach millimoles 06 ascorbic acid : 100 -3.75 = 96.25 mmol millimoles of sodium ascorbate = 150+ 3.75 = 153.